# Fair Resource Allocation in Multi-Task Learning

### v1

@@ 5 Empirical Results
We conduct extensive experiments under both supervised learning and reinforcement learning settings to demonstrate the effectiveness of our proposed method. Full experimental details and more empirical studies can be found in Appendix A.
@@ 5.1 Supervised Learning
Image-Level Classification. CelebA [55] is a large-scale face attributes dataset, containing over 200K celebrity images. Each image is annotated with 40 attributes, such as smiling, wavy hair, mustache, etc. We can consider the dataset as an image-level 40-task MTL classification problem, with each task predicting the presence of a specific attribute. This setting assesses the capability of MTL methods in handling a large number of tasks. We follow the experiment setup in [28]. We employ a network containing a 9-layer convolutional neural network (CNN) as the backbone and a specific linear layer for each task. We train our method for 15 epochs, using Adam optimizer with learning rate 3e-4. The batch size is 256.
@@
Regression. QM9 [56] is a widely-used benchmark in graph neural networks. It comprises over 130k organic molecules, which are organized as graphs with annotated node and edge features. The goal of predicting 11 properties with different measurement scales is to see if MTL methods can effectively balance the variations present across these tasks. Following [26; 28], we use the example provided in Pytorch Geometric [57], and use 110k molecules for training, 10k for validation, and the rest 10k for testing. We train our method for 300 epochs with a batch size of 120. The initial learning rate is 1e-3, and a scheduler is used to reduce the learning rate once the improvement of validation stagnates.
@@
Dense Prediction. NYU-v2 [58] contains 1449 densely annotated images that have been collected from video sequences of various indoor scenes. It involves one pixel-level classification task and two pixel-level regression tasks, which correspond to 13-class semantic segmentation, depth estimation, and surface normal prediction, respectively. Similarly, Cityscapes [59] contains 5000 street-scene images with two tasks: 7-class semantic segmentation and depth estimation. This scenario evaluates the effectiveness of MTL methods in tackling complex situations. We follow [21; 26; 28] and adopt the backbone of MTAN [37], which adds task-specific attention modules on SegNet [60]. We train our method for 200 epochs with batch size 2 for NYU-v2 and 8 for Cityscapes. The learning rate is 1e-4 for the first 100 epochs, then decayed by half for the rest.
Method CelebA QM9 MR normal-$\downarrow$ $\Delta m\%\downarrow$ MR normal-$\downarrow$ $\Delta m\%\downarrow$ LS 6.53 4.15 8.18 177.6 SI 8.00 7.20 4.82 77.8 RLW 5.40 1.46 9.55 203.8 DWA 7.23 3.20 7.82 175.3 UW 6.00 3.23 6.18 108.0 MGDA 11.05 14.85 7.73 120.5 PCGrad 6.98 3.17 6.36 125.7 CAGrad 6.53 2.48 7.18 112.8 IMTL-G 4.95 0.84 6.09 77.2 Nash-MTL 5.38 2.84 3.64 62.0 FAMO 5.03 1.21 4.73 58.5 FairGrad 4.95 0.37 3.82 57.9
@@
Evaluation. For image-level classification and regression, we compare our FairGrad with Linear Scalarization (LS) which minimizes the sum of task losses, Scale-Invariant (SI) which minimizes the sum of logarithmic losses, Random Loss Weighting (RLW) [61], Dynamic Weigh Average (DWA) [37], Uncertainty weighting (UW) [43], MGDA [23], PCGrad [20], CAGrad [21], IMTL-G [62], Nash-MTL [26], and FAMO [28]. For dense prediction, we also compare with GradDrop [63], and MoCo [29]. We consider two metrics to represent the overall performance of the MTL method $m$. (1) $\mathbf{\Delta m\%}$, the average per-task performance drop against the single-task (STL) baseline $b$:
@@
Method Segmentation Depth Surface Normal MR normal-$\downarrow$ $\Delta m\%\downarrow$ mIoU normal-$\uparrow$ Pix Acc normal-$\uparrow$ Abs Err normal-$\downarrow$ Rel Err normal-$\downarrow$ Angle Distance normal-$\downarrow$ Within $t^{\circ}$ normal-$\uparrow$ Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 10.67 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 9.44 4.39 RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 13.11 7.78 DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 9.44 3.57 UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 9.22 4.05 MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 7.11 1.38 PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 9.78 3.97 GradDrop 39.39 65.12 0.5455 0.2279 27.48 22.96 23.38 49.44 62.87 8.78 3.58 CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 5.78 0.20 IMTL-G 39.35 65.60 0.5426 0.2256 26.02 21.19 26.20 53.13 66.24 5.11 -0.76 MoCo 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 5.44 0.16 Nash-MTL 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 3.11 -4.04 FAMO 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 4.44 -4.10 FairGrad 39.74 66.01 0.5377 0.2236 24.84 19.60 29.26 56.58 69.16 2.67 -4.66
Method Segmentation Depth MR normal-$\downarrow$ $\Delta m\%\downarrow$ mIoU normal-$\uparrow$ Pix Acc normal-$\uparrow$ Abs Err normal-$\downarrow$ Rel Err normal-$\downarrow$ STL 74.01 93.16 0.0125 27.77 LS 75.18 93.49 0.0155 46.77 8.50 22.60 SI 70.95 91.73 0.0161 33.83 10.50 14.11 RLW 74.57 93.41 0.0158 47.79 10.75 24.38 DWA 75.24 93.52 0.0160 44.37 8.50 21.45 UW 72.02 92.85 0.0140 30.13 6.75 5.89 MGDA 68.84 91.54 0.0309 33.50 11.00 44.14 PCGrad 75.13 93.48 0.0154 42.07 8.50 18.29 GradDrop 75.27 93.53 0.0157 47.54 8.00 23.73 CAGrad 75.16 93.48 0.0141 37.60 7.00 11.64 IMTL-G 75.33 93.49 0.0135 38.41 5.50 11.10 MoCo 75.42 93.55 0.0149 34.19 4.50 9.90 Nash-MTL 75.41 93.66 0.0129 35.02 3.25 6.82 FAMO 74.54 93.29 0.0145 32.59 7.25 8.13 FairGrad 75.72 93.68 0.0134 32.25 1.50 5.18
@@
Furthermore, there are some other interesting findings from the results presented in Table 2. LS performs poorly in the surface normal prediction (SNP) task compared to the other two tasks. This is because LS does not take the fairness among tasks into consideration, and hence the gradient of the SNP task is dominated by the others. On the contrary, MGDA [23] obtains the best performance in the SNP task among all three tasks by enforcing the max-min fairness. Meanwhile, the performance of Nash-MTL [26], which embodies proportional fairness, is more balanced across all tasks. As a comparison, our FairGrad can find a better-balanced solution than LS and MGDA, while placing greater emphasis on the challenging SNP tasks than Nash-MTL.
@@ 5.2 Reinforcement Learning
We further evaluate our method on the MT10, a benchmark including 10 robotic manipulation tasks from the MetaWorld environment [64], where the objective is to learn one policy that generalizes to different tasks such as pick and place, open door, etc. We follow [21; 26; 28] and adopt Soft Actor-Critic (SAC) [65] as the underlying algorithm. We implement with MTRL codebase [66] and train our method for 2 million steps with a batch size of 1280.
Method success rate (mean ± stderr) STL 0.90 $\pm$ 0.03 MTL SAC 0.49 $\pm$ 0.07 MTL SAC + TE 0.54 $\pm$ 0.05 MH SAC 0.61 $\pm$ 0.04 PCGrad 0.72 $\pm$ 0.02 CAGrad 0.83 $\pm$ 0.05 MoCo 0.75 $\pm$ 0.05 ${}^{\star}$ 0.91 $\pm$ 0.03 Nash-MTL (reproduced) 0.80 $\pm$ 0.13 FAMO 0.83 $\pm$ 0.05 FairGrad 0.84 $\pm$ 0.07
Table 4: Results on MT10 benchmark. Average over 10 random seeds. ${}^{\star}$ denotes the result reported in the original paper [26]. While Nash-MTL (reproduced) denotes the reproduced result in [28].
Evaluation. We compare our FairGrad with Multi-task SAC (MTL SAC) [64], Multi-task SAC with task encoder (MTL SAC + TE) [64], Multi-headed SAC (MH SAC) [64], PCGrad [20], CAGrad [21], MoCo [29], Nash-MTL [26], and FAMO [28].
Results. The results are shown in Table 4. Each method is evaluated once every 10,000 steps, and the best average success rate over 10 random seeds throughout the entire training course is reported. We could not reproduce the MTRL result in the original paper of Nash-MTL exactly, and hence we adopt the reproduced result of Nash-MTL in [28]. It is evident that our method performs competitively when compared to other methods.
@@ 5.3 Effect of Different Fairness Criteria
Also note that although the best $\Delta m\%$ reported in Table 5 is better than that reported in Table 2 and Table 3, their results w.r.t. MR are worse than those in Table 2 and Table 3. This is because an improved $\Delta m\%$ may result in a lower rank for certain tasks, causing a significant degradation in the average rank.
@@ 5.4 Discussion on Practical Implementation
How do we choose $\alpha$. We first search with $\alpha\in[1,2,5,10]$ and evaluate which choice is better. Then we narrow down the search space and continue to execute a grid search with a step size of 0.1 until we determine an appropriate value.
Supervised Learning. For experiments on QM9, NYU-v2, and Cityscapes, we implement our method based on the codes released by [26]. For experiments on CelebA, our implementation is based on the codes provided by [28], consistent with all the baselines presented in Table 1.
Reinforcement Learning. We find it time-consuming to solve the constrained nonlinear least square problem discussed in Section 4 under the reinforcement learning setting. Therefore, we use SGD to approximately solve the problem and accelerate the training process. Specifically, we use SGD optimizer with a learning rate of 0.1 and train 20 epochs.
Method Segmentation Depth Surface Normal $\Delta m\%\downarrow$ mIoU normal-$\uparrow$ Pix Acc normal-$\uparrow$ Abs Err normal-$\downarrow$ Rel Err normal-$\downarrow$ Angle Distance normal-$\downarrow$ Within $t^{\circ}$ normal-$\uparrow$ Mean Median 11.25 22.5 30 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 5.59 Fair-LS 38.64 64.96 0.5422 0.2255 27.14 22.64 24.05 50.14 63.53 2.85 RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 7.78 Fair-RLW 37.29 63.58 0.5481 0.2263 27.67 23.33 23.38 48.72 62.12 5.00 DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 3.57 Fair-DWA 39.03 65.18 0.5404 0.2266 27.20 22.63 24.31 50.14 63.45 2.65 UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 4.05 Fair-UW 38.51 64.56 0.5423 0.2274 27.23 22.92 23.62 49.52 63.23 3.56 MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 1.38 Fair-MGDA 35.91 63.19 0.5646 0.2260 24.75 19.24 30.04 57.30 69.55 -3.26 PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 3.97 Fair-PCGrad 39.26 65.08 0.5257 0.2177 26.88 22.26 24.74 50.85 64.18 1.23 CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 0.20 Fair-CAGrad 39.32 65.36 0.5290 0.2221 25.50 20.32 28.06 54.94 67.65 -2.91

### v2

@@ 5 Empirical Results
We first use a toy example to elaborate how FairGrad balances the tasks by incorporating different fairness criteria. Then we conduct extensive experiments under both supervised learning and reinforcement learning settings to demonstrate the effectiveness of our proposed method. Full experimental details and more empirical studies can be found in Appendix A.
5.1 Toy Example
We adopt the 2-task toy example introduced in (Navon et al., 2022), where the objectives of the 2 tasks, denoted as $L_{1}$ and $L_{2}$, have different scales. More details are provided in Section A.1. We select 5 starting points and illustrate the optimization trajectories of FairGrad with different fairness criteria in Figure 1.
Obviously, without fairness (Linear Scalarization), the algorithm may not converge to a Pareto stationary point. However, in other cases where fairness is involved, the algorithm can converge. Furthermore, in the experiment setting, objective $L_{2}$ exhibits a larger scale than objective $L_{1}$, resulting in a larger gradient magnitude. If there is no fairness, task 2 will dominate the optimization process. When the algorithm converges, it will always converge to a stationary point where $L_{2}$ is smaller than $L_{1}$, as shown in Figure 1. On the other hand, with max-min fairness, the least fortunate task will be prioritized. The algorithm tends to converge to a stationary point with a smaller $L_{1}$. Proportional fairness and MPD fairness will lead to a more balanced solution.
Method CelebA QM9 MR $\downarrow$ $\Delta m\%\downarrow$ MR $\downarrow$ $\Delta m\%\downarrow$ LS 6.53 4.15 8.18 177.6 SI 8.00 7.20 4.82 77.8 RLW 5.40 1.46 9.55 203.8 DWA 7.23 3.20 7.82 175.3 UW 6.00 3.23 6.18 108.0 MGDA 11.05 14.85 7.73 120.5 PCGrad 6.98 3.17 6.36 125.7 CAGrad 6.53 2.48 7.18 112.8 IMTL-G 4.95 0.84 6.09 77.2 Nash-MTL 5.38 2.84 3.64 62.0 FAMO 5.03 1.21 4.73 58.5 FairGrad 4.95 0.37 3.82 57.9
Method Segmentation Depth Surface Normal MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 STL 38.30 63.76 0.6754 0.2780 25.01 19.21 30.14 57.20 69.15 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 10.67 5.59 SI 38.45 64.27 0.5354 0.2201 27.60 23.37 22.53 48.57 62.32 9.44 4.39 RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 13.11 7.78 DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 9.44 3.57 UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 9.22 4.05 MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 7.11 1.38 PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 9.78 3.97 GradDrop 39.39 65.12 0.5455 0.2279 27.48 22.96 23.38 49.44 62.87 8.78 3.58 CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 5.78 0.20 IMTL-G 39.35 65.60 0.5426 0.2256 26.02 21.19 26.20 53.13 66.24 5.11 -0.76 MoCo 40.30 66.07 0.5575 0.2135 26.67 21.83 25.61 51.78 64.85 5.44 0.16 Nash-MTL 40.13 65.93 0.5261 0.2171 25.26 20.08 28.40 55.47 68.15 3.11 -4.04 FAMO 38.88 64.90 0.5474 0.2194 25.06 19.57 29.21 56.61 68.98 4.44 -4.10 FairGrad 39.74 66.01 0.5377 0.2236 24.84 19.60 29.26 56.58 69.16 2.67 -4.66
@@ 5.2 Supervised Learning
Image-Level Classification. CelebA (Liu et al., 2015) is a large-scale face attributes dataset, containing over 200K celebrity images. Each image is annotated with 40 attributes, such as smiling, wavy hair, mustache, etc. We can consider the dataset as an image-level 40-task MTL classification problem, with each task predicting the presence of a specific attribute. This setting assesses the capability of MTL methods in handling a large number of tasks. We follow the experiment setup in (Liu et al., 2023). We employ a network containing a 9-layer convolutional neural network (CNN) as the backbone and a specific linear layer for each task. We train our method for 15 epochs, using Adam optimizer with learning rate 3e-4. The batch size is 256.
@@
Regression. QM9 (Ramakrishnan et al., 2014) is a widely-used benchmark in graph neural networks. It comprises over 130k organic molecules, which are organized as graphs with annotated node and edge features. The goal of predicting 11 properties with different measurement scales is to see if MTL methods can effectively balance the variations present across these tasks. Following (Navon et al., 2022; Liu et al., 2023), we use the example provided in Pytorch Geometric (Fey & Lenssen, 2019), and use 110k molecules for training, 10k for validation, and the rest 10k for testing. We train our method for 300 epochs with a batch size of 120. The initial learning rate is 1e-3, and a scheduler is used to reduce the learning rate once the improvement of validation stagnates.
@@
Dense Prediction. NYU-v2 (Silberman et al., 2012) contains 1449 densely annotated images that have been collected from video sequences of various indoor scenes. It involves one pixel-level classification task and two pixel-level regression tasks, which correspond to 13-class semantic segmentation, depth estimation, and surface normal prediction, respectively. Similarly, Cityscapes (Cordts et al., 2016) contains 5000 street-scene images with two tasks: 7-class semantic segmentation and depth estimation. This scenario evaluates the effectiveness of MTL methods in tackling complex situations. We follow (Liu et al., 2021; Navon et al., 2022; Liu et al., 2023) and adopt the backbone of MTAN (Liu et al., 2019), which adds task-specific attention modules on SegNet (Badrinarayanan et al., 2017). We train our method for 200 epochs with batch size 2 for NYU-v2 and 8 for Cityscapes. The learning rate is 1e-4 for the first 100 epochs, then decayed by half for the rest.
@@
Evaluation. For image-level classification and regression, we compare our FairGrad with Linear Scalarization (LS) which minimizes the sum of task losses, Scale-Invariant (SI) which minimizes the sum of logarithmic losses, Random Loss Weighting (RLW) (Lin et al., 2021), Dynamic Weigh Average (DWA) (Liu et al., 2019), Uncertainty weighting (UW) (Kendall et al., 2018), MGDA (Sener & Koltun, 2018), PCGrad (Yu et al., 2020b), CAGrad (Liu et al., 2021), IMTL-G (Liu et al., 2020), Nash-MTL (Navon et al., 2022), and FAMO (Liu et al., 2023). For dense prediction, we also compare with GradDrop (Chen et al., 2020), and MoCo (Fernando et al., 2022). We consider two metrics to represent the overall performance of the MTL method $m$. (1) $\mathbf{\Delta m\%}$, the average per-task performance drop against the single-task (STL) baseline $b$:
@@
Method Segmentation Depth MR $\downarrow$ $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ STL 74.01 93.16 0.0125 27.77 LS 75.18 93.49 0.0155 46.77 8.50 22.60 SI 70.95 91.73 0.0161 33.83 10.50 14.11 RLW 74.57 93.41 0.0158 47.79 10.75 24.38 DWA 75.24 93.52 0.0160 44.37 8.50 21.45 UW 72.02 92.85 0.0140 30.13 6.75 5.89 MGDA 68.84 91.54 0.0309 33.50 11.00 44.14 PCGrad 75.13 93.48 0.0154 42.07 8.50 18.29 GradDrop 75.27 93.53 0.0157 47.54 8.00 23.73 CAGrad 75.16 93.48 0.0141 37.60 7.00 11.64 IMTL-G 75.33 93.49 0.0135 38.41 5.50 11.10 MoCo 75.42 93.55 0.0149 34.19 4.50 9.90 Nash-MTL 75.41 93.66 0.0129 35.02 3.25 6.82 FAMO 74.54 93.29 0.0145 32.59 7.25 8.13 FairGrad 75.72 93.68 0.0134 32.25 1.50 5.18
@@
Furthermore, there are some other interesting findings from the results presented in Table 2. LS performs poorly in the surface normal prediction (SNP) task compared to the other two tasks. This is because LS does not take the fairness among tasks into consideration, and hence the gradient of the SNP task is dominated by the others. On the contrary, MGDA (Sener & Koltun, 2018) obtains the best performance in the SNP task among all three tasks by enforcing the max-min fairness. Meanwhile, the performance of Nash-MTL (Navon et al., 2022), which embodies proportional fairness, is more balanced across all tasks. As a comparison, our FairGrad can find a more balanced solution than LS and MGDA, while placing greater emphasis on the challenging SNP tasks than Nash-MTL.
@@ 5.3 Reinforcement Learning
We further evaluate our method on the MT10, a benchmark including 10 robotic manipulation tasks from the MetaWorld environment (Yu et al., 2020c), where the objective is to learn one policy that generalizes to different tasks such as pick and place, open door, etc. We follow (Liu et al., 2021; Navon et al., 2022; Liu et al., 2023) and adopt Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the underlying algorithm. We implement with MTRL codebase (Shagun Sodhani, 2021) and train our method for 2 million steps with a batch size of 1280.
Table 4: Results on MT10 benchmark. Average over 10 random seeds. Nash-MTL denotes the result reported in the original paper (Navon et al., 2022). While Nash-MTL (reproduced) denotes the reproduced result in (Liu et al., 2023).
Method success rate (mean ± stderr) STL 0.90 $\pm$ 0.03 MTL SAC 0.49 $\pm$ 0.07 MTL SAC + TE 0.54 $\pm$ 0.05 MH SAC 0.61 $\pm$ 0.04 PCGrad 0.72 $\pm$ 0.02 CAGrad 0.83 $\pm$ 0.05 MoCo 0.75 $\pm$ 0.05 Nash-MTL 0.91 $\pm$ 0.03 Nash-MTL (reproduced) 0.80 $\pm$ 0.13 FAMO 0.83 $\pm$ 0.05 FairGrad 0.84 $\pm$ 0.07
Evaluation. We compare our FairGrad with Multi-task SAC (MTL SAC) (Yu et al., 2020c), Multi-task SAC with task encoder (MTL SAC + TE) (Yu et al., 2020c), Multi-headed SAC (MH SAC) (Yu et al., 2020c), PCGrad (Yu et al., 2020b), CAGrad (Liu et al., 2021), MoCo (Fernando et al., 2022), Nash-MTL (Navon et al., 2022), and FAMO (Liu et al., 2023).
Results. The results are shown in Table 4. Each method is evaluated once every 10,000 steps, and the best average success rate over 10 random seeds throughout the entire training course is reported. We could not reproduce the MTRL result in the original paper of Nash-MTL exactly, and hence we adopt the reproduced result of Nash-MTL in (Liu et al., 2023). It is evident that our method performs competitively when compared to other methods.
@@ 5.4 Effect of Different Fairness Criteria
Also note that although the best $\Delta m\%$ reported in Table 5 is better than that reported in Table 2 and Table 3, their results w.r.t. MR are worse than those in Table 2 and Table 3. This is because an improved $\Delta m\%$ may result in a lower rank for certain tasks, causing a significant degradation in the average rank. See Section A.3 for more details.
@@ 5.5 Discussion on Practical Implementation
Supervised Learning. For experiments on QM9, NYU-v2, and Cityscapes, we implement our method based on the codes released by (Navon et al., 2022). For experiments on CelebA, our implementation is based on the codes provided by (Liu et al., 2023), consistent with all the baselines presented in Table 1.
Reinforcement Learning. We find it time-consuming to solve the constrained nonlinear least square problem discussed in Section 4 under the reinforcement learning setting. Therefore, we use SGD to approximately solve the problem and accelerate the training process. Specifically, we use SGD optimizer with a learning rate of 0.1, momentum of 0.5, and train 20 epochs.
Choice of $\alpha$. We first search with $\alpha\in[1,2,5,10]$ and evaluate which choice is better. Then we narrow down the search space and continue to execute a grid search with a step size of 0.1 until we determine an appropriate value.
In practice, the choice depends on the specific needs or preferences. If there are no requirements for fairness and tasks with larger gradients are allowed to finish first, we can simply set $\alpha=0$ to allow for quick training. If tasks with struggling progress are prioritized (e.g., in some meta-learning setups, some harder-to-train tasks may play more important roles in deciding final test accuracy), then the max-min fairness (with a larger $\alpha$) is desired. From our observation, if aiming to achieve the most balanced overall performance, MPD fairness with $\alpha=2$ is preferable. After fairness criteria are selected, some slight finetuning on $\alpha$ can also be conducted to further improve the overall accuracy.
@@
Method Segmentation Depth Surface Normal $\Delta m\%\downarrow$ mIoU $\uparrow$ Pix Acc $\uparrow$ Abs Err $\downarrow$ Rel Err $\downarrow$ Angle Distance $\downarrow$ Within $t^{\circ}$uparrow Mean Median 11.25 22.5 30 LS 39.29 65.33 0.5493 0.2263 28.15 23.96 22.09 47.50 61.08 5.59 Fair-LS 38.64 64.96 0.5422 0.2255 27.14 22.64 24.05 50.14 63.53 2.85 RLW 37.17 63.77 0.5759 0.2410 28.27 24.18 22.26 47.05 60.62 7.78 Fair-RLW 37.29 63.58 0.5481 0.2263 27.67 23.33 23.38 48.72 62.12 5.00 DWA 39.11 65.31 0.5510 0.2285 27.61 23.18 24.17 50.18 62.39 3.57 Fair-DWA 39.03 65.18 0.5404 0.2266 27.20 22.63 24.31 50.14 63.45 2.65 UW 36.87 63.17 0.5446 0.2260 27.04 22.61 23.54 49.05 63.65 4.05 Fair-UW 38.51 64.56 0.5423 0.2274 27.23 22.92 23.62 49.52 63.23 3.56 MGDA 30.47 59.90 0.6070 0.2555 24.88 19.45 29.18 56.88 69.36 1.38 Fair-MGDA 35.91 63.19 0.5646 0.2260 24.75 19.24 30.04 57.30 69.55 -3.26 PCGrad 38.06 64.64 0.5550 0.2325 27.41 22.80 23.86 49.83 63.14 3.97 Fair-PCGrad 39.26 65.08 0.5257 0.2177 26.88 22.26 24.74 50.85 64.18 1.23 CAGrad 39.79 65.49 0.5486 0.2250 26.31 21.58 25.61 52.36 65.58 0.20 Fair-CAGrad 39.32 65.36 0.5290 0.2221 25.50 20.32 28.06 54.94 67.65 -2.91